### v1

@@ 5.3 Manual evaluation
We selected 20 papers, ten from each dataset (note that we use the test subset in case of CSPubSum) for human evaluation. A human annotator working in the field of NLP was asked to choose the most appropriate title among the generated ones. It is found that for CSPubSum dataset, in $80\%$ of the cases, the title generated by PEGASUS-large is most preferred, while for LREC-COLING-2024, in 50% cases, the output of PEGASUS-large wins while in 40% cases, the title generated by ChatGPT-3.5 is considered the best.
@@ 6 Demo
A demo of the application is hosted at https://title-generation-researchpapers.onrender.com/ where the user can input an abstract into a text box, select a suitable fine-tuned language model, and the maximum token count for the title, and obtain a title generated by the model. We have found that if the generated title is incomplete, increasing the token count (to say, 25 or 30) generally produces a correct and complete title.
Figure 4: Graphical user interface of our pre-trained language model-based title generation application.
@@ 7 Conclusion
We applied six different pre-trained/LLM models to generate titles of research papers. The fine-tuned PEGASUS-large model has achieved the best performance on the CSPubSum dataset as well the LREC-COLING-2024 dataset. Our extensive experiments suggest that pre-trained and large language models can generate engaging and suitable titles for research papers. The success of PEGASUS-large suggests that lighter pre-trained language models, fine-tuned on domain-specific datasets, may be satisfactory and even more suitable for this task than large language models with billions of parameters and substantial computational requirements. The high scores achieved by the models on LREC-COLING-2024 dataset on which they are not fine-tuned indicates that fine-tuning on a related scholarly dataset suffices to generate acceptable performance. Although the generated titles are generally of high quality, authors may further refine them to better suit their papers. This tool can be particularly useful for non-native English speakers and novice researchers. In future, we would like to analyze in more detail the differences between author-assigned and AI-generated titles and explore techniques to assess the quality of the abstractive nature of the outputs from LLMs.

### v2

@@ 5.3 Manual evaluation
We selected twenty papers, ten from each dataset (for CSPubSum, they were sampled from the test subset) for human evaluation. A human annotator working in the field of NLP was asked to choose the most appropriate title among the ones generated by the six models. It was found that for CSPubSum, in $80\%$ of the cases, the titles generated by PEGASUS-large were the most preferred. For the LREC-COLING-2024 dataset, PEGASUS-large was the winning model in 50% of cases, while the titles generated by GPT-3.5-turbo were considered the best in 40% of cases.
5.4 Creative titles with LLMs
We explored whether ChatGPT can generate creative titles for papers. In particular, we prompted ChatGPT-4o to generate titles in specific styles – namely, catchy, rhyming, funny, and comprehensive – one style at a time for ten papers from CSPubSum (test set) and ten from LREC-COLING-2024. An example is presented in Fig. 4, which displays titles in the above four styles along with the ‘plain’ title generated by GPT-3.5-turbo (using the prompt in Section 4.2) and the author-written one. We depict the similarity of these titles with the author-written ones in Table 5 (ignore the last column). Catchy, comprehensive, and plain titles exhibit similar similarity scores. However, rhyming and funny titles generally share very few words (sometimes with zero bi-gram overlap) and only moderate semantic similarity with the author-written titles. This is not surprising, as these styles often introduce novel words not present in the original titles, and tend to sacrifice technical terms to adhere to the stylistic constraints.
Author-written title: “A Tool for Determining Distances and Overlaps between Multimodal Annotations”
Catchy: “Analyzing ELAN Annotations: A Tool for Comparing Multimodal Overlaps in Corpus Analysis”
Rhyming: “Comparative Insights in ELAN’s Light: A Tool for Annotations That Delivers Clear Sight”
Funny: “Annotation Adventures: Finding Overlaps and Other Mischief!”
Comprehensive: “Tool for Analyzing ELAN Annotations: Comparing Speech and Gesture Overlap in Multimodal Corpora”
Plain: “Tool for Comparing ELAN Annotations in Multimodal Corpus Analysis”
Figure 4: Titles generated by ChatGPT for an abstract in LREC-COLING-2024. Paper taken from https://aclanthology.org/2024.lrec-main.150/.
Table 5: Evaluation of ChatGPT-generated creative titles.
Dataset Style ROUGE-1 ROUGE-2 ROUGE-L METEOR MoverScore BERTScore SciBERTScore Avg. Rating CSPubSum Catchy 45.81 18.44 38.83 45.55 35.43 88.00 75.77 4 Rhyming 13.20 0 11.37 10.37 12.32 83.47 61.02 3.3 Funny 18.57 28.18 14.95 15.09 14.76 83.78 59.53 2.8 Compreh. 48.03 29.32 40.88 47.93 44.21 88.16 76.78 4.4 Plain 47.51 25.03 42.09 44.99 40.74 88.15 77 3.4 LREC-COLING-2024 Catchy 47.82 24.47 41.60 45.61 42.84 90.07 75.77 3.8 Rhyming 20.48 02.65 16.94 16.64 20.72 84.37 61.02 3.1 Funny 15.71 0.0 11.36 10.11 11.47 84.05 59.53 3.6 Compreh. 48.22 21.88 41.12 48.83 45.71 90.4 76.78 4.1 Plain 43.83 19.45 39.76 39.42 42.48 90.88 77.00 3.7
We selected nine human annotators, all undergraduate or graduate students in computer science, to rate the ChatGPT-generated titles on a scale of 1-5. The average ratings are reported in the last column of Table 5. Comprehensive titles have scored the highest. Author-written titles for both datasets received an average rating of 4. When asked to choose the best title for each paper, the annotators preferred ChatGPT-generated titles to the author-written ones for nine LREC-COLING-2024 papers and seven CSPubSum papers, mostly favoring the comprehensive or catchy titles.
@@ 6 Demo
A demo of the application is publicly hosted. As shown in Fig. 5, a user can input an abstract into a text box, select a suitable fine-tuned language model and the maximum token count for the title, and obtain a title generated by the model. We have found that if the generated title is incomplete, increasing the token count (to say, 25 or 30) generally produces a correct and complete title.
Figure 5: Graphical user interface of our title generation application.
@@ 7 Conclusion
We applied six different pre-trained language models to generate titles of research papers. The fine-tuned PEGASUS-large model has achieved the best performance on the CSPubSum dataset as well the LREC-COLING-2024 dataset. Our extensive experiments indicate that these language models can generate engaging and suitable titles for research papers. The success of PEGASUS-large suggests that lighter pre-trained language models, fine-tuned on domain-specific datasets, may be satisfactory and even more suitable for this task than large language models with billions of parameters and substantial computational requirements. The high scores achieved by the models on LREC-COLING-2024 dataset, on which they were not fine-tuned, indicates that fine-tuning on a related scholarly dataset suffices to generate acceptable performance. Although the generated titles are generally of high quality, authors might refine them further to better suit their papers. This tool can be particularly useful for non-native English speakers and novice researchers. In future, we would like to conduct a more detailed analysis of the difference between author-written and AI-generated titles, as well as explore better techniques for evaluating the abstractive quality of the outputs from LLMs.